OpenStudy (anonymous):

Solve the following system of equations by using the substitution method: t = 4-2s, (1) and t + 2s = 6, (2)

OpenStudy (anonymous):

ha ha

OpenStudy (anonymous):

what do the (1) and the (2) stand for? t?? or s??

OpenStudy (anonymous):

I am not sure if those commas are supposed to be there or not. The instructor put a comma before the (1) on the first equation, but he didn't on the second equation.

OpenStudy (anonymous):

Is your instructor Bon Jovi?

OpenStudy (anonymous):

LOL, no why?

OpenStudy (anonymous):

I'm assuming that the (1) is going to take the place of the "s", and not the "t". but really, it's not specified...

OpenStudy (anonymous):

The exact way he worded it was t = 4 - 2s, (1) and for the second equation he put t + 2s = 6 (2)

OpenStudy (anonymous):

Does it say to solve for t? or for s....

OpenStudy (anonymous):

So I think there is either an omission or a typo in there someplace.

OpenStudy (anonymous):

Yes, solve using the substitution method.

OpenStudy (anonymous):

he doesn't say whether to solve for t or s.

OpenStudy (anonymous):

But we have to solve for either T.... Or we have to solve for S

OpenStudy (anonymous):

then he is a silly goose.

OpenStudy (anonymous):

Well, he does a lot of copying and pasting his questions, so he may have just not worded it correctly.

OpenStudy (anonymous):

here...

OpenStudy (anonymous):

t = 4 - 2(s) , (1) put the one in for s..."substitute" and get t = 4 - 2(1)

OpenStudy (anonymous):

2*1 = 2 so its now..... t = 4-2 = 2

OpenStudy (anonymous):

t=2 for the first...

OpenStudy (anonymous):

second one...

OpenStudy (anonymous):

t + 2s = 6 , (2)

OpenStudy (anonymous):

t + 2(2) = 6..........t + 4 = 6...........t = 2 (after subtracting 4 from both sides

OpenStudy (anonymous):

So basically both equations have the same answer?

OpenStudy (anonymous):

yea...

OpenStudy (anonymous):

okay, I will post this to his question and ask him if he is sure he asked it correctly. LOL thank you Ryan.

OpenStudy (anonymous):

they only have the same answer because there were different values for s that we put in namely 1, and 2.. they aren't the same equation though... later

OpenStudy (anonymous):

right. thank you.

OpenStudy (saifoo.khan):

heyy!